# BS-LDM: Effective Bone Suppression in High-Resolution Chest X-Ray Images with Conditional Latent Diffusion Models

### v3

@@ Abstract
Lung diseases represent a significant global health challenge, with Chest X-Ray (CXR) being a key diagnostic tool due to their accessibility and affordability. Nonetheless, the detection of pulmonary lesions is often hindered by overlapping bone structures in CXR images, leading to potential misdiagnoses. To address this issue, we developed an end-to-end framework called BS-LDM, designed to effectively suppress bone in high-resolution CXR images. This framework is based on conditional latent diffusion models and incorporates a multi-level hybrid loss-constrained vector-quantized generative adversarial network which is crafted for perceptual compression, ensuring the preservation of details. To further enhance the framework’s performance, we introduce offset noise and a temporal adaptive thresholding strategy. These additions help minimize discrepancies in generating low-frequency information, thereby improving the clarity of the generated soft tissue images. Additionally, we have compiled a high-quality bone suppression dataset named SZCH-X-Rays. This dataset includes 818 pairs of high-resolution CXR and dual-energy subtraction soft tissue images collected from a partner hospital. Moreover, we processed 241 data pairs from the JSRT dataset into negative images, which are more commonly used in clinical practice. Our comprehensive experimental and clinical evaluations reveal that BS-LDM excels in bone suppression, underscoring its significant clinical value.
@@ 1 Introduction
Lung disease remains a leading cause of high morbidity and mortality worldwide [1, 2]. Chest X-Ray (CXR) is the primary imaging technique for evaluating conditions such as inflammation, tuberculosis, and lung masses owing to its accessibility, cost-effectiveness, and low radiation exposure. However, even skilled radiologists may miss lesions in CXR images. This challenge primarily arises from bone structures overlapping with the lung region, which can obscure crucial details necessary for accurate detection and diagnosis. Research indicates that between 82% and 95% of undetected lung cancers are concealed by these bone structures [3]. Consequently, bone suppression techniques in CXR images could greatly aid radiologists and enhance the performance of computer-aided pulmonary lesion detection [4, 5].
@@
With advancements in neural networks, deep learning-based algorithms have increasingly been used for bone suppression in CXR images, generating soft tissue images effectively. These methods typically approach bone suppression as either an end-to-end image denoising task or focus on predicting the bone residual. Generally, end-to-end image denoising yields better bone suppression, but it may compromise detail preservation due to model limitations. Conversely, the bone residual prediction approach often produces impressive results in the generation of lesion and texture, although it struggles to completely remove bones. Yang et al. [11] developed a cascaded multi-scale Convolutional Neural Network (CNN) trained within the gradient domain of CXR images for this purpose. While the model demonstrated impressive performance, it failed to maintain high levels of perceptual and structural integrity. Similarly, Gusarev et al. [12] treated bones as noise, using a combination of autoencoder and deep CNN features to remove bone structures, though this led to image blurring. Inspired by Generative Adversarial Networks (GANs) [13], Zhou et al. [14] introduced a Multi-scale Conditional Adversarial Network (MCA-Net) aimed at producing soft tissue images while preserving key anatomical structures. To address the issue of inconsistent background intensity in gradient-based methods, Chen et al. [15] introduced a cascaded CNN into the wavelet domain. Rajaraman et al. [16] developed the ResNet-BS model for bone suppression in CXR images, validating its effectiveness through subsequent analytic tasks. Additionally, Liu et al. [17] proposed a bone suppression technique for lateral CXR images utilizing data rectification and distillation learning.
@@
To apply diffusion models to bone suppression, Chen et al. [21] introduced BS-Diff, which involves developing a conditional diffusion model in pixel space, enhanced with a Variational Auto-Encoder (VAE) as an improvement module. The cascade architecture used in BS-Diff helps mitigate some low-frequency discrepancies typical of diffusion models. However, the model still faces challenges in achieving adequate resolution and fine detail, alongside high computational demands. Moreover, utilizing the cascade architecture at the same resolution can lead to partial information loss while incurring additional computational costs. Consequently, the challenges associated with diffusion models in bone suppression remain only partially addressed.
@@
Figure 1: Overview of the proposed BS-LDM: (a) The training process of BS-LDM, where CXR and noised soft tissue data in the latent space are transmitted to the noise estimator for offset noise prediction and L2 loss calculation; (b) The training process of ML-VQGAN, where a multi-level hybrid loss-constrained VQGAN is used to construct a latent space by training the reconstruction of CXR and soft tissue images, using a codebook to represent the discrete features of the images; (c) The sampling process of BS-LDM, where the latent variables obtained after each denoising step are clipped using a temporal adaptive thresholding strategy for the sake of contrast stability.
@@
Recently, Latent Diffusion Models (LDMs) [22] have emerged as an extension of diffusion models, operating in a lower-dimensional space to enable the synthesis of high-resolution images with reduced computational costs. By leveraging latent space, these models significantly decrease the number of parameters in the noise estimator network. Building on this innovation, we propose an end-to-end LDM-based framework for high-resolution bone suppression, called BS-LDM, which treats the task as CXR-guided image denoising to achieve effective bone suppression. To enhance detail preservation, we design a multi-level hybrid loss-constrained Vector-Quantized Generative Adversarial Network (VQGAN) [23], known as ML-VQGAN, for perceptual compression of high-resolution images. Additionally, to address discrepancies in low-frequency information generation, we introduce offset noise [24] into the noise addition process to correct luminance offset during image generation. We also propose a temporal adaptive thresholding strategy to resolve pixel saturation issues during the reverse process. To assess the clinical relevance of our framework in aiding lung disease diagnosis, we conducted a clinical evaluation focusing on image quality and diagnostic utility. The results demonstrated excellent image quality scores and significant diagnostic improvements, highlighting the clinical value of our approach. Furthermore, we compiled a high-quality bone suppression dataset named SZCH-X-Rays, featuring high-resolution paired CXR and DES soft tissue images from 818 patients, collected in collaboration with our partner hospital. We also processed 241 pairs of CXR and DES soft tissue images from the JSRT dataset, the largest open-source dataset for bone suppression, by performing operations like inversion and contrast adjustment to convert these images into negative formats more frequently used in clinical settings. Our primary contributions are summarized as follows:
  • We introduce an end-to-end LDM-based framework for high-resolution bone suppression, named BS-LDM. It utilizes a multi-level hybrid loss-constrained VQGAN for effective perceptual compression. This framework consistently generates soft tissue images with high levels of bone suppression while preserving fine details and critical lung lesions.
@@
  • To enhance the quality of generated images, we incorporate offset noise and a temporal adaptive thresholding strategy. These innovations help minimize discrepancies in low-frequency information, thereby improving the interpretability of the soft tissue images.
  • We have compiled a comprehensive bone suppression dataset, SZCH-X-Rays, which includes 818 pairs of high-resolution CXR and DES soft tissue images from our partner hospital. Additionally, we processed 241 pairs of images from the JSRT dataset into negative formats more commonly used in clinical settings.
@@
  • Our clinical evaluation focused on image quality and diagnostic utility. The results demonstrated excellent image quality scores and substantial diagnostic improvements, underscoring the clinical significance of our work.
@@ 2.1 Overall Architecture
This section presents the proposed BS-LDM framework in detail. It is an end-to-end framework designed for effective bone suppression in high-resolution CXR images using a conditional LDM, as depicted in Fig. 1. In this framework, we developed a multi-level hybrid loss-constrained VQGAN for perceptual compression to ensure good detail preservation. Additionally, we incorporated offset noise and a temporal adaptive thresholding strategy to address discrepancies in low-frequency information generation, enhancing the overall quality of generation. A detailed description of each component of BS-LDM is provided in the following subsections.
@@ 2.2 Conditional Latent Diffusion Models
LDM-based approaches, such as Stable Diffusion [22], typically operate through a two-stage generative process. First, they map data to a low-dimensional latent space, and then sample from this space to generate the final output. Unlike traditional diffusion models like Denoising Diffusion Probabilistic Models (DDPMs) [18], LDMs significantly reduce computational demands for generating high-resolution images by encoding data in a compressed latent space. Specifically, for an image $x\in\mathbb{R}^{C\times H\times W}$, the encoder $E$ transforms $x$ into a latent variable $z=E(x)$. In these models, a forward process, denoted as $q_{\theta}(z_{t}|z_{0})$, is defined using a Markovian chain that gradually adds Gaussian noise to the input data $z_{0}$:
@@
where $\epsilon_{g}$ is a noise map sampled from a Gaussian distribution, and $\bar{\alpha}_{t}:=\prod_{s=0}^{t}\alpha_{s}$. Here, $\alpha_{t}=1-\beta_{t}$ is a differentiable function of timestep $t$ determined by the denoising sampler. The diffusion training loss is formulated as:
$\mathcal{L}\left(\epsilon_{\theta}\right)=\sum_{t=1}^{T}\mathbb{E}_{z_{0},\epsilon_{g}}\left[\left\|\epsilon_{\theta}\left(\sqrt{\bar{\alpha}_{t}}z_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{g}\right)-\epsilon_{g}\right\|_{2}^{2}\right].$ (2)
In the context of conditional LDMs, with a given image condition $\widetilde{z}$, the diffusion training loss at timestep $t$ can be re-written as:
@@
where $\epsilon_{\theta}$ represents the predicted noise at timestep $t$ by the noise estimator network with parameters $\theta$, and $\sigma_{t}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}$ is the variance of posterior Gaussian distribution $p_{\theta}\left(x_{0}\right)$.
In this work, we employ a U-Net [25] with multi-resolution attention [26] as the noise estimator network. The conditional guidance of CXR images in BS-LDM is achieved through the channel concatenation of $z_{t}$ and $\widetilde{z}$, as illustrated in Fig. 1 (a).
@@ 2.3 Multi-Level Hybrid Loss-Constrained VQGAN
In this formulation, $\mathcal{L}_{Recon}$ is a combination of multiple reconstruction losses at different levels. The weight coefficients $\lambda_{Qua}$, $\lambda_{L1}$, $\lambda_{Per}$, and $\lambda_{Adv}$ adjust the influence of the respective components $\mathcal{L}_{Qua}$, $\mathcal{L}_{L1}$, $\mathcal{L}_{Per}$, and $\mathcal{L}_{Adv}$ on the total loss.
@@ 2.3.2 Quantization Loss
In ML-VQGAN, the concept of quantization loss, $\mathcal{L}_{Qua}$, as described by Tudosiu et al. [30], is used to optimize the codebook $\mathcal{Z}$. The mathematical expression for this loss is:
@@ 2.4 Offset Noise
In the forward process of LDMs, each pixel gradually accumulates a small amount of independently and identically distributed (i.i.d.) Gaussian noise at each step. LDMs are trained to predict this noise after it has disrupted real images. However, we observe that low-frequency features are more resistant to Gaussian noise, as illustrated in Fig. 2. To understand this phenomenon, we calculated the power spectral densities of soft tissue images from SZCH-X-Rays, corresponding latent variables, and Gaussian noise at different frequencies, as shown in Fig. 3.
Our analysis revealed that the power spectral densities of the low-frequency components in both real images and latent variables are higher compared to mid- and high-frequency components. In contrast, Gaussian noise maintains balanced power spectral densities across all frequency bands. This imbalance may explain why low-frequency components are less susceptible to corruption during the noise addition process in diffusion models and LDMs. As a result, low-frequency information is not entirely erased, leading to a noise signature that differs from standard Gaussian noise. Consequently, during the reverse process, which begins with Gaussian noise, there is a lack of consistency with the forward process. This inconsistency causes low-frequency information in generated images to resemble Gaussian noise rather than real images, potentially biasing the mean value of generated soft tissue images towards Gaussian noise. This bias can result in luminance discrepancies, reducing the interpretability of the images.
To resolve this issue, we introduce offset noise $\epsilon$ during the forward process, as depicted in Fig. 1 (a). This involves adding extra bias noise onto the standard noise addition procedure, as shown in Fig. 4 and formalized in (11):
@@ 2.5 Temporal Adaptive Thresholding Strategy
Although thresholding is typically absent in latent diffusion models [32], we observed that without it, these models are prone to pixel saturation, resulting in overly contrasted outputs. Static thresholding offers some numerical stability and reduces pixel saturation, but in our task, it severely restricted image variance, creating a marked contrast discrepancy between generated and real images. Similarly, while dynamic thresholding provides more flexibility, it still fails to adequately generate soft tissue images with clear interpretability.
To address these issues and align pixel intensities of generated soft tissue images with real ones, we propose a temporal adaptive thresholding strategy. This method dynamically adjusts the thresholding range $[-s,s]$ according to the current sampling step $t$. For the thresholding operation $\tau$, we utilize a linear function to define the relationship between the threshold $s$ and current sampling timestep $t$, as shown in (12):
@@
Figure 5: Visualization of soft tissue images on SZCH-X-Rays and JSRT datasets produced by different methods.

### v5

@@ Abstract
Lung diseases represent a significant global health challenge, with Chest X-Ray (CXR) being a key diagnostic tool due to its accessibility and affordability. Nonetheless, the detection of pulmonary lesions is often hindered by overlapping bone structures in CXR images, leading to potential misdiagnoses. To address this issue, we develop an end-to-end framework called BS-LDM, designed to effectively suppress bone in high-resolution CXR images. This framework is based on conditional latent diffusion models and incorporates a multi-level hybrid loss-constrained vector-quantized generative adversarial network which is crafted for perceptual compression, ensuring the preservation of details. To further enhance the framework’s performance, we utilize offset noise in the forward process, and a temporal adaptive thresholding strategy in the reverse process. These additions help minimize discrepancies in generating low-frequency information of soft tissue images. Additionally, we have compiled a high-quality bone suppression dataset named SZCH-X-Rays. This dataset includes 818 pairs of high-resolution CXR and soft tissue images collected from our partner hospital. Moreover, we processed 241 data pairs from the JSRT dataset into negative images, which are more commonly used in clinical practice. Our comprehensive experiments and downstream evaluations reveal that BS-LDM excels in bone suppression, underscoring its clinical value. Our code is available at https://github.com/diaoquesang/BS-LDM.
@@ 1 Introduction
Lung disease remains a leading cause of high morbidity and mortality worldwide [1, 2]. Chest X-Ray (CXR) is the primary imaging technique for evaluating conditions such as inflammation, tuberculosis, and lung masses owing to its accessibility, cost-effectiveness, and low radiation exposure. However, even skilled radiologists may miss lesions in CXR images. This challenge primarily arises from bone structures overlapping with the lung region, which can obscure crucial details necessary for accurate detection and diagnosis. Research indicates that between 82% and 95% of undetected lung cancers are concealed by these bone structures [3]. Consequently, bone suppression techniques in CXR images could greatly aid radiologists [4], and enhance the performance of computer-aided pulmonary lesion detection [5].
@@
With advancements in neural networks, deep learning-based algorithms have increasingly been used for bone suppression in CXR images, generating soft tissue images effectively. These methods typically approach bone suppression as either an end-to-end image denoising task or focus on predicting the bone residual. Generally, end-to-end image denoising yields better bone suppression, but it may compromise detail preservation due to model limitations. Conversely, the bone residual prediction approach often produces impressive results in the generation of lesion and texture, although it struggles to completely remove bones. Yang et al. [11] presented a cascaded multi-scale Convolutional Neural Network (CNN) trained within the gradient domain of CXR images for this purpose. While the model demonstrated impressive performance, it failed to maintain high levels of perceptual and structural integrity. Similarly, Gusarev et al. [12] treated bones as noise, using a combination of autoencoder and deep CNN features to remove bone structures, though this led to image blurring. Inspired by Generative Adversarial Networks (GANs) [13], Zhou et al. [14] introduced a Multi-scale Conditional Adversarial Network (MCA-Net) aimed at producing soft tissue images while preserving key anatomical structures. To address the issue of inconsistent background intensity in gradient-based methods, Chen et al. [15] employed a cascaded CNN into the wavelet domain. Rajaraman et al. [16] developed the ResNet-BS model for bone suppression in CXR images, validating its effectiveness through subsequent analytic tasks. Liu et al. [17] proposed a bone suppression technique for lateral CXR images utilizing data rectification and distillation learning. Rani et al. [18] utilized a Pix2Pix-based conditional Generative Adversarial Network (cGAN) for effective bone suppression. Additionally, Singh et al. [19] advanced this work with the Multiscale and Dual Attention Generative Adversarial Network (MDA-GAN), which incorporates a multiscale and dual attention mechanism into a conditional generative adversarial network, effectively removing bone structures while preserving background and chest X-ray details. Furthermore, Schiller et al. [20] provided the xU-NetFullSharp architecture with multi-level and multi-directional connections to remove bone shadows in CXR images.
@@
To apply diffusion models to bone suppression, Chen et al. [24] introduced BS-Diff, which involves developing a conditional diffusion model in pixel space, enhanced with an Auto-Encoder as an enhancement module. The cascade architecture used in BS-Diff helps mitigate some low-frequency discrepancies typical of diffusion models. However, the model still faces challenges in achieving adequate resolution and fine detail, alongside high computational demands. Moreover, utilizing the cascade architecture at the same resolution may lead to partial information loss while incurring additional computational costs. Consequently, the challenges associated with diffusion models in bone suppression remain only partially addressed.
@@
Figure 1: Overview of the proposed BS-LDM: (a) The training process of BS-LDM, where CXR and noised soft tissue data in the latent space are transmitted to the noise estimator for offset noise prediction and L2 loss calculation; (b) The training process of ML-VQGAN, where a multi-level hybrid loss function is used to construct a latent space by reconstruction training while preserving texture details; (c) The sampling process of BS-LDM, where the latent variables obtained after each denoising step are clipped using a temporal adaptive thresholding strategy for the sake of contrast stability.
@@
Recently, Latent Diffusion Models (LDMs) [25] have emerged as an extension of diffusion models, operating in a lower-dimensional space to enable the synthesis of high-resolution images with reduced computational costs. By leveraging latent space, these models significantly decrease the number of parameters in the noise estimator network. Building on this innovation, we propose an end-to-end LDM-based framework for high-resolution bone suppression, called BS-LDM, which treats the task as CXR-guided image denoising to achieve effective bone suppression. To enhance detail preservation, we design a Multi-Level hybrid loss-constrained Vector-Quantized Generative Adversarial Network (ML-VQGAN), for perceptual compression of high-resolution images. Additionally, to address discrepancies in low-frequency information generation by compensating low-frequency disruption in the forward process, we introduce offset noise [26] into the noise addition process to correct luminance offset during image generation. Additionally, we propose a temporal adaptive thresholding strategy to resolve pixel saturation issues during the reverse process. To assess the clinical relevance of our framework in aiding lung disease diagnosis, we conducted a clinical evaluation focusing on image quality and diagnostic utility. The results demonstrated excellent image quality scores and significant diagnostic improvements, highlighting the clinical value of our approach. Our automated downstream evaluation results further demonstrate the potential of our BS-LDM model for enhancing the performance of computer-assisted lung lesion detection. Furthermore, we compiled a high-quality bone suppression dataset named SZCH-X-Rays, featuring high-resolution paired CXR and DES soft tissue images from 818 patients, collected in collaboration with our partner hospital. We also processed 241 pairs of CXR and DES soft tissue images from the JSRT dataset, the largest open-source dataset for bone suppression, by performing operations like inversion and contrast adjustment to convert these images into negative formats more frequently used in clinical settings. Our primary contributions are summarized as follows:
  • We propose BS-LDM, an LDM-based framework for high-resolution bone suppression in CXR images, and introduce ML-VQGAN for effective perceptual compression and detail retention. To the best of our knowledge, this is the first study to train and evaluate LDMs for the bone suppression task.
@@
  • To enhance the quality of generated images, we incorporate offset noise in the forward process, and temporal adaptive thresholding in the reverse process. These strategies help minimize discrepancies in low-frequency information, thereby improving the interpretability of the soft tissue images.
@@
  • Our comprehensive experiments, as well as clinical and automated downstream evaluations demonstrated excellent image quality and substantial diagnostic improvements, underscoring the clinical significance of our work.
@@ 2.1 Overall Architecture
This section presents the proposed BS-LDM framework in detail. It is an end-to-end framework designed for effective bone suppression in high-resolution CXR images using a conditional LDM, as depicted in Fig. 1. In this framework, we developed ML-VQGAN for perceptual compression to ensure good detail preservation. Additionally, we incorporated offset noise in the forward process and temporal adaptive thresholding in the reverse process to address discrepancies in low-frequency information generation, enhancing the overall quality of generation. A detailed description of each component of BS-LDM is provided in the following subsections.
@@ 2.2 Conditional Latent Diffusion Models
LDM-based approaches [27, 28, 29], typically operate through a two-stage generative process. First, they map data to a low-dimensional latent space, and then sample from this space to generate the final output. Unlike traditional diffusion models [21], LDMs significantly reduce computational demands for high-resolution image synthesis by encoding data in a compressed latent space. Specifically, for an image $x\in\mathbb{R}^{C\times H\times W}$, the encoder $E$ transforms $x$ into a latent variable $z=E(x)$. In the context of conditional LDMs, a forward process, denoted as $q_{\theta}(z_{t}|z_{0})$, is defined as a process gradually adding Gaussian noise to the input data $z_{0}$:
@@
where $\epsilon_{g}$ is a noise map sampled from a Gaussian distribution, and $\bar{\alpha}_{t}:=\prod_{s=0}^{t}\alpha_{s}$. Here, $\alpha_{t}=1-\beta_{t}$ is a differentiable function of timestep $t$ determined by the denoising sampler. The diffusion training loss with a given image condition $\widetilde{z}$ at timestep $t$ is formulated as:
@@
where $\epsilon_{\theta}$ represents the predicted noise by the noise estimator network with parameters $\theta$ at timestep $t$, and $\sigma_{t}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}$ is the variance of posterior Gaussian distribution $p_{\theta}\left(x_{0}\right)$.
In this work, we employ a U-Net with multi-resolution attention as the noise estimator network. The conditional guidance of CXR images in BS-LDM is achieved through channel concatenation of $z_{t}$ and $\widetilde{z}$, as illustrated in Fig. 1 (a).
@@ 2.3 Multi-Level Hybrid Loss-Constrained VQGAN
In this formulation, $\mathcal{L}_{Recon}$ is a combination of different reconstruction losses at multiple levels. The weight coefficients $\lambda_{Qua}$, $\lambda_{L1}$, $\lambda_{Per}$, and $\lambda_{Adv}$ adjust the influence of the respective components $\mathcal{L}_{Qua}$, $\mathcal{L}_{L1}$, $\mathcal{L}_{Per}$, and $\mathcal{L}_{Adv}$ on the total loss.
@@ 2.3.2 Quantization Loss
In ML-VQGAN, the concept of quantization loss, $\mathcal{L}_{Qua}$, is used to optimize the codebook $\mathcal{Z}$. The mathematical expression for this loss is:
@@ 2.4 Offset Noise
In the forward process of LDMs, each pixel gradually accumulates a small amount of independently and identically distributed (i.i.d.) Gaussian noise at each step. LDMs are trained to predict this noise after it has disrupted real images. Nervertheless, when implementing bone suppression using LDMs, we observed a notable discrepancy in their performance: while LDMs effectively generated high-frequency textures, their handling of low-frequency information was suboptimal. To investigate this issue, we conducted a spectral decomposition analysis of the latent variables of soft tissue images using the Discrete Fourier Transform. This analysis revealed differences in how conventional Gaussian noise injection impacts frequency bands during the generative process. As illustrated in Fig. 2, low-frequency features exhibit greater resistance to Gaussian noise compared to high-frequency features. This disparity affects the training dynamics of the noise estimator network, leading to uneven estimation capabilities across frequency components. Specifically, the reduced ability to denoise low-frequency features forces LDMs to retain certain elements of the initialization point, namely, Gaussian noise, as low-frequency information during the reverse generation process. Consequently, this manifests as a shift in the mean value of the generated images towards the mean of the Gaussian noise. To further explore the underlying causes of this phenomenon, we calculated the power spectral densities of soft tissue images from the SZCH-X-Rays dataset, alongside their corresponding latent variables and Gaussian noise across various frequencies. The results, presented in Fig. 3, provide deeper insights into the interplay between frequency components and Gaussian noise during the LDM generative process. The results demonstrate that the power spectral density of Gaussian noise is relatively uniform across all frequencies, with slight fluctuations around a constant value. This indicates that the Gaussian noise has a consistent intensity at different frequencies. In contrast, the power spectral density of soft tissue images and their corresponding latent variables exhibits a distinct pattern. The analysis showed that the power spectral density is higher in the low-frequency region and decreases progressively in the mid- to high-frequency regions. This indicates a relatively greater intensity of low-frequency information compared to higher-frequency components. This difference can be attributed to the inherent characteristics of soft tissue images and the nature of Gaussian noise. Soft tissue images contain rich low-frequency information which represent the overall shape and intensity variations of the tissues. These low-frequency features are crucial for the visual perception and diagnostic value of the images.
To address this issue, we apply offset noise $\epsilon$ into the forward process to compensate the disturbance of low-frequency information, as illustrated in Fig. 1(a). This approach involves incorporating additional zero-frequency bias noise into the standard noise injection procedure, as demonstrated in Fig. 4 and mathematically defined in Equation (10):
@@ 2.5 Temporal Adaptive Thresholding Strategy
Although thresholding is typically absent in latent diffusion models [34], we observed that without it, these models are prone to pixel saturation, resulting in overly contrasted outputs. While static thresholding provides numerical stability and helps mitigate pixel saturation, it significantly limits image variance in our task. This limitation results in noticeable contrast differences between the generated images and real ones, as shown in Fig. 7. Similarly, while dynamic thresholding provides more flexibility, it still fails to adequately generate soft tissue images with clear interpretability. In LDMs, the distribution of the latent variable $z_{t}$ changes over time as $t$ progresses. Applying static or dynamic thresholding based solely on pixel statistics can introduce biased contrast, as it fails to account for the temporal dynamics of the latent variable.
To address these issues and align pixel intensities of generated soft tissue images with real ones, we propose a temporal adaptive thresholding strategy. This method dynamically adjusts the thresholding range $[-s,s]$ according to the current sampling step $t$. For the thresholding operation $\tau$, we utilize a linear function to define the relationship between the threshold $s$ and current sampling timestep $t$, as shown in Equation (11):